OpenStudy (anonymous):

Write an equation in slope-intercept form of the line satisfying the given conditions. Through (-9, 5), perpendicular to -3x + 2y = 21 I got y=2/3x-7 is that right? > _ <'

OpenStudy (anonymous):

Also in the equation x/2 + 8x/9 =x/18 , x would just be 0, right?

OpenStudy (anonymous):

And Write a slope-intercept equation of the line whose graph is described. Parallel to the graph of 6x + 7y = 5; y-intercept (0, 1) I got y= 6/7x + 1 is that right? ''0 ___ o

OpenStudy (anonymous):

nevermind e - e'